OpenStudy (anonymous):

Radians: What is the exact value of tan 5/3 pi

OpenStudy (anonymous):

\[\tan 5\div3 \pi \]

OpenStudy (anonymous):

can you please show the working out :)

OpenStudy (zehanz):

The tangent has period pi, this means you can add or subtract pi from the angle as many times as you like, the tangent remains the same, or: \[\tan \frac{ 5 }{ 3 }\pi = \tan \left( \frac{ 5 }{ 3 }\pi -2\pi \right) = \tan \left( -\frac{ 1 }{ 3 } \pi\right) \] Now you can place the minus sign in front of the tan, because it is an uneven function. This means that for every x: tan(-x) = -tanx, so: \[\tan \left( -\frac{ 1 }{ 3 }\pi \right) = -\tan \left( \frac{ 1 }{ 3 } \pi\right)\] Now the hardest part ;) : Aren't you supposed to learn by heart som special values of sin, cos and tan? Well, if you are, you now remember that\[\tan \left( \frac{ 1 }{ 3 }\pi \right) = \sqrt{3}\] So the answer is \[-\sqrt{3}\] Hope this helps! ZeHanz

OpenStudy (anonymous):

You are the best thank you